OpenStudy (anonymous):

Solve: 0 = 4x2 − 8x − 5. x=-1/2,x=5/2 x=1/2,x=-5/2 x=1/2,x=-5/2 x=1/2,x=5/2

OpenStudy (anonymous):

Do you know how to factor by grouping?

OpenStudy (anonymous):

dont i have to multiply the 4 and 8 and the substrac the 2 by the x?

OpenStudy (anonymous):

No. Find factors of -20 that add up to -8.

OpenStudy (anonymous):

-28

OpenStudy (anonymous):

How is -28 a factor of -20 and how does it add to -8?

OpenStudy (anonymous):

You're not supposed to add them.

OpenStudy (anonymous):

0 = 4x^2 - 8x - 5 0 = 4x^2 - 10x + 2x - 5 0 = 2x(2x - 5) + (2x - 5) 0 = (2x + 1)(2x - 5) Can you solve for x from here?